Investors require a 17% rate of return on Levine Company's stock (i.e., rs = 17%).

  1. What is its value if the previous dividend was D0 = $3.00 and investors expect dividends to grow at a constant annual rate of (1) -2%, (2) 0%, (3) 4%, or (4) 12%? Do not round intermediate calculations. Round your answers to two decimal places.

    (1) $

    (2) $

    (3) $

    (4) $
Answer #1

The question is solved using the dividend discount model.

1.Information provided:

Current dividend= $3.00

Required rate of return= 17%

Growth rate= -2%

Price of the stock today=D1/(r-g)

where:

D1=next dividend payment

r=interest rate

g=firm’s expected growth rate

Price of stock today= $3*(1 -0.02)/ 0.17 + 0.02

                                = $2.94/ 0.19

                                = $15.47.

Therefore, the value of the stock at -2% growth rate is $15.47.

2.Information provided:

Current dividend= $3.00

Required rate of return= 17%

Growth rate= 0%

The question is solved using the dividend discount model.

Price of the stock today=D1/(r-g)

where:

D1=next dividend payment

r=interest rate

g=firm’s expected growth rate

Price of stock today= $3*(1+ 0)/ 0.17 – 0

                                = $3/ 0.17

  = $17.65.

Therefore, the value of the stock at 0% growth rate is $17.65.

3.Information provided:

Current dividend= $3.00

Required rate of return= 17%

Growth rate= 4%

The question is solved using the dividend discount model.

Price of the stock today=D1/(r-g)

where:

D1=next dividend payment

r=interest rate

g=firm’s expected growth rate

Price of the stock today= $3*(1 + 0.04)/ 0.17- 0.04

                                      = 3.12/ 0.13          

  =$24

Therefore, the value of the stock at 4% growth rate is $24.

4.Information provided:

Current dividend= $3.00

Required rate of return= 17%

Growth rate= 12%

The question is solved using the dividend discount model.

Price of the stock today=D1/(r-g)

where:

D1=next dividend payment

r=interest rate

g=firm’s expected growth rate

Price of the stock today= $3*(1+0.12)/ 0.17 – 0.12

                                     = $3.36/ 0.05

                                     = $67.20.

Therefore, the value of the stock at 12% growth rate is $67.20.

In case of any further queries, kindly comment on the solution.